# The Dual of Quantifier Elimination: Boolean Elimination over $\mathbb{C}$ and $\mathbb{R}$

### v1

@@ 1. Introduction
Section 7 turns to other base fields, and shows that over $\mathbb{R}$, any Boolean combination of equations and inequations can be represented with a single $\exists$ quantifying a polynomial equality. Section 8 constructs the $\exists^{d}$ and $\forall\exists$ polynomials for Boolean combinations over $\mathbb{R}$ that may involve inequalities such as $u>0$. Section 9 notes prior results over $\mathbb{R}$, and section 10 concludes with a brief exploration of forms over $\mathbb{Q}$.
@@ 2. Notation
The constructions below are uniform in the dimension of $\mathbb{C}^{n}$; the quantified variables $a,b,r,s,y,z$ always range over scalars. Throughout, an inequation of the form $u(x)\neq 0$ refers simply to non-vanishing in the ambient field, whether $\mathbb{C}$ or $\mathbb{R}$. The order conditions $u>0$ or $u\geq 0$ appear only in sections 8 and 9.
@@ 3. Context
The third key ingredient is basic quantifier elimination over an algebraically closed field, where we need only the elimination of a single existentially quantified variable:
@@ 7. Extending the results on equations from $\mathbb{C}$ to $\mathbb{R}$
A simple existential form is available for $\mathbb{R}$, or any other formally real field such as $\mathbb{Q}$, so long as we retain the restriction to formulas with equations and inequations.
@@ 8. Normalizing to $\exists^{d}$ and $\forall\exists$ over $\mathbb{R}$
For inequalities in $\mathbb{R}$, it is typical to also consider order conditions such as $u>0$. We can in fact convert Boolean combinations of such inequalities over $\mathbb{R}$ to both $\exists^{d}$ and $\forall\exists$ forms. We state theorems and constructions analogous to the main theorem here over $\mathbb{C}$, whose main difference from the complex cases is the use of results in Lemma 3.1 with >> instead of $\neq$. We omit detailed proofs, especially since we will later cite other results establishing an $\exists$ form for these formulas.
@@
9. Shorter $\exists$ forms over $\mathbb{R}$
David Marker called our attention to a result of Motzkin providing an $\exists$ normal form over $\mathbb{R}$. This result was later proved by Pecker with better degrees, as follows:
@@ 9. Shorter $\exists$ forms over $\mathbb{R}$
Let $\varphi(x)$ be a finite combination of polynomial equalities and order-inequalities in conjunctive normal form. Then there is a polynomial $p(r,x)\in\mathbb{R}[r,x]$, of degree $4$ in $r$, such that, for all $x\in\mathbb{R}^{n}$,
@@
Simplifying some of Pecker’s polynomials gives:
$(y>0\wedge z>0)$ $\iff(\exists r)(1+y-r^{2}yz)^{2}=y$
$(x>0\wedge y>0\wedge z>0)$ $\iff(\exists r)[(1+x+y-r^{2}xyz)^{2}-x-y]^{2}=xy$
@@ 10. Normal forms over $\mathbb{Q}$
We can also reduce the $4$’s to $3$’s in these signatures, by using a fact from number theory and a corollary reformulation:
@@
$\forall u\in\mathbb{Q}\Big[u>0\iff(\exists v_{1},v_{2},v_{3}\in\mathbb{Q})\,[1-u(v_{1}^{2}+v_{2}^{2}+v_{3}^{2})][1-2u(v_{1}^{2}+v_{2}^{2}+v_{3}^{2})]=0\Big].$
@@
The results here show that the Boolean structure of constructible or semialgebraic sets can be uniformly represented with a single polynomial equation under either an $\exists\forall$ or a $\forall\exists$ prefix for $\mathbb{C}$, or even under a simple $\exists$ prefix for $\mathbb{R}$, and that those results are optimal. We hope it will be useful to have those results conveniently assembled here. These results also point to the possibility of Boolean elimination elsewhere, and the examples of both $\mathbb{R}$ and $\mathbb{Q}$ show that finding short forms and proving their optimality may require substantial results.

### v2

@@ 1. Introduction
Section 7 turns to other base fields, and shows that any Boolean combination of equations and inequations over $\mathbb{R}$ ($u\neq 0$) can be represented with a single $\exists$ and a single equality. Section 8 constructs the $\exists^{d}$ and $\forall\exists$ polynomials for Boolean combinations over $\mathbb{R}$ that may involve inequalities ($u>0$). Sections 9 and 10 explore prior results over $\mathbb{R}$ and forms over $\mathbb{Q}$. Section 11 concludes with directions for further research.
@@ 2. Notation
The constructions below are uniform in the dimension of $\mathbb{C}^{n}$, and where we need to refer to a vector there, we use $x$. The quantified variables $a,b,r,s,v,w,y,z$ always range over scalars, whether in $\mathbb{C}$, $\mathbb{R}$, or $\mathbb{Q}$.
@@ 3. Context
The third key ingredient is basic quantifier elimination over an algebraically closed field, where we need only the elimination of a single existentially quantified equation:
@@
$(\forall z\in\mathbb{C})\,(q(z)\neq 0)$ $\iff q_{0}\neq 0\wedge q_{1}=0\wedge\cdots\wedge q_{d}=0$
@@ 7. Extending the results on equations from $\mathbb{C}$ to $\mathbb{R}$
A simple existential form is available for $\mathbb{R}$, or any other formally real field such as $\mathbb{Q}$, so long as we retain the restriction to formulas with equations and inequations. We switch notation for quantified variables from $a,b$ to $r,s$ to reflect the change from $\mathbb{C}$ to $\mathbb{R}$.
@@ 8. Normalizing to $\exists^{d}$ and $\forall\exists$ over $\mathbb{R}$
For inequalities in $\mathbb{R}$, it is typical to also consider order conditions such as $u>0$, and indeed Boolean combinations of such inequalities over $\mathbb{R}$ can be converted to both $\exists^{d}$ and $\forall\exists$ forms. We state theorems and constructions here analogous to the main theorem over $\mathbb{C}$, using the real-specific results in Lemma 3.1 with >> instead of $\neq$. We omit detailed proofs, especially since we will later cite other results establishing an $\exists$ form for these formulas.
@@
9. Single-quantifier $\exists$ forms over $\mathbb{R}$
The previous sections gave explicit $\exists^{d}$ and $\forall\exists$ forms for Boolean combinations of inequalities over $\mathbb{R}$. David Marker called our attention to a result of Motzkin showing that in fact a single existential quantifier always suffices. This result was later sharpened by Pecker, who achieved the single quantifier via polynomials with degrees larger than the polynomials above, but much smaller than Motzkin’s. We may write his result in our terms as an $\exists$ normal form:
@@ 9. Single-quantifier $\exists$ forms over $\mathbb{R}$
Let $\varphi(x)$ be a finite combination of polynomial equalities and $f$ strict inequalities in conjunctive normal form. Then there is a polynomial $p(r,x)\in\mathbb{R}[r,x]$, of degree $2^{f}$ in $r$, such that, for all $x\in\mathbb{R}^{n}$,
@@
As an example, consider the polynomial produced in Pecker’s first corollary:
$(y>0\wedge z>0)\iff(\exists r)(r^{2}yz-y^{2}z-1)^{2}-y^{2}z=0$
We may simplify this to
$(y>0\wedge z>0)\iff(\exists r)(r^{2}yz-z-1)^{2}=z$
Solving the simplified equation for $y$ gives $y=(z+1\pm\sqrt{z})/r^{2}z$, which makes it easier to see that it has roots exactly when $y$ and $z$ are positive.
Figure 1. Graph for a simple version of Pecker’s polynomial
As $r$ takes all possible values, graphs like the above sweep through the entire first quadrant.
@@ 10. Normal forms over $\mathbb{Q}$
We can also reduce the $4$’s to $3$’s in these signatures, by using a fact from number theory and a corollary reformulation. We use $v,w$ for quantified variables to reflect the switch from $\mathbb{C}$ and $\mathbb{R}$ to $\mathbb{Q}$:
@@
$u\in\mathbb{Q}\wedge u>0\iff(\exists v_{1},v_{2},v_{3}\in\mathbb{Q})\,[1-u(v_{1}^{2}+v_{2}^{2}+v_{3}^{2})][1-2u(v_{1}^{2}+v_{2}^{2}+v_{3}^{2})]=0\Big$.
@@
11. Conclusion
The results here show that the Boolean structure of constructible or semialgebraic sets can be uniformly represented with a single polynomial equation under either an $\exists\forall$ or a $\forall\exists$ prefix for $\mathbb{C}$, or even under a simple $\exists$ prefix for $\mathbb{R}$, and that those quantifier patterns are optimal. We hope it will be useful to have those results conveniently assembled here.
These results also lead to questions about various extensions of them.
  • Can the $\exists$ results about $\mathbb{R}$ be simplified or optimized for smaller degrees, and can the $\exists^{3}$ results about $\mathbb{Q}$ be reduced to fewer quantifiers?
  • With quantified input formulas, can every universally quantified formula be normalized to $\forall\exists$ form with a single equation, and every existentially quantified formula to $\exists\forall$ form?
  • What variants of these constructions are available when the Boolean combinations may be countable as in $\mathcal{L}_{\omega_{1},\omega}$ rather than finite, and the language may be expanded beyond polynomials to include exponentials or other univariate entire functions?
These questions point beyond elementary methods toward a deeper study of Boolean elimination.